# Computational Design and Fabrication of Modular Robots with Untethered Control

### v1

@@ Abstract
Natural organisms use distributed actuation via their musculoskeletal systems to adapt their gait for traversing diverse terrains or to morph their bodies to perform varied tasks. A longstanding challenge in the field of robotics is to mimic this extensive adaptability and range of motion. This has led humans to develop various soft robotic systems that emulate natural organisms. However, such systems are generally optimized for a single functionality, lack the ability to change form or function on demand, or are often tethered to bulky control systems. To address these challenges, we present our framework for designing and controlling robots that mimic nature’s blueprint by utilizing distributed actuation. We propose a novel building block that combines 3D-printed bones with liquid crystal elastomer (LCE) muscles as lightweight actuators and enables the modular assembly of musculoskeletal robots. We developed LCE rods that contract in response to infrared radiation, thereby achieving local and untethered control over the distributed network of bones, which in turn results in global deformation of the robot. Furthermore, to capitalize on the extensive design space, we develop two computational tools: one to optimize the robot’s skeletal graph, enabling multiple target deformations, and another to co-optimize the skeletal designs and control gaits to achieve target locomotion. We validate our system by building several robots that show complex shape morphing, varying control schemes, and adaptability to their environment. Our system integrates advances in modular material building, untethered and distributed control, and computational design to introduce a new generation of robots that brings us closer to the capabilities of living organisms.
1 Introduction
Figure 1: Our building block (a) is a 3D-printed bone equipped with two LCE muscles (b) that act as untethered actuators when heated. This block can undergo reversible deformation, to either preferentially bend or shrink (c). We use the building block as an edge in the skeletal graph (d) to design our modular robot. The skeletal graph is optimized for different terrains and is used to generate the CAD design (e) from which the robot is 3D printed (f) and the muscles (in green) are attached. An infrared radiation controlled with a galvo-scanner (g) enables the untethered and active control of the robots (h).
@@ 1 Introduction
Natural organisms have developed remarkable mobility to traverse challenging terrains and perform complex tasks. Their ability to readily adapt to their environment lies in their structural design; instead of specialized appendages, they leverage general building blocks shaped by evolution. Consider vertebrates, arguably the most successful animal group; their diverse species possess evolutionary-optimized skeletons tailored to their specific habitats. Further enhancing this design, a distributed network of muscles empowers these skeletons for intricate motions. This synergy creates a powerful and versatile blueprint that enables vertebrates to thrive in a wide range of environments and tasks.
@@
Figure 2: Active control of tethered soft robots (a,b) [1, 2] has showcased rich performance for shape morphing and locomotion tasks. In contrast, untethered control achieved through passive actuation via heat (c) [3] , light (d) [4] and magnetism (e) [5] remains limited in performance. Untethered robots with active control combine the best of both approaches and have been used for space exploration (f) [6] and development of insect-scale flapping wing vehicle [7]. From a design perspective, modular robots have shown promise by combining basic building blocks to perform shape-morphing (h) [8] or locomotion [9] tasks. Computational design tools allow for structural optimization of different shape morphing structures that are passively controlled using heat (j) [10], elastic deformation (k) [11] or inflation [12]. Modular robots created using computational design tools promise significant potential as demonstrated by the conceptual robots in [m,n,o] [13, 14, 15]. Our system (p) uniquely integrates these different domains enabling computational design of modular robots with both active and untethered control to perform shape morphing and locomotion tasks.
@@
This contrasts with the capabilities of human-designed robots that mimic natural organisms—soft robots. Soft robots have gained traction in the robotics community due to their ability to interact safely with humans or delicate environments. They are typically made of soft elastic material, where a single spatial location can generally undergo two modes of deformation: it can either bend and/or shrink. These modes of deformation have inspired many researchers to assemble individual blocks into structures that can be passively actuated into target shapes. [11, 16] introduced an auxetic building block that exhibits a negative poison ratio to passively transform a flat-fabricated configuration into target 3D shapes. [17, 18] utilized building blocks composed of elastic beams, rigid plates, and hinge joints to design mechanically deployable structures. [12] introduced inflatable fusing curves as building blocks in which the flat design is encoded to passively inflate to a target 3D shape. [10, 19] demonstrated how shrinkable materials can serve as building blocks, passively actuated with hot water to achieve the desired 3D shapes. Similar building blocks were used by [20] to print a heat-shrinkable base sheet that self-folds into target origami objects. However, without active control of each individual unit, these systems remain limited in their ability to support general-purpose soft robotic applications.
@@
On the flip side, researchers proposed robotic materials embedded with actively controlled actuators. [1, 21] designed robotic surfaces with high spatio-temporal control to perform shape morphing and object manipulation. [2, 22, 23] designed shape morphing robots capable of performing locomotion tasks across a variety of terrains. Active control of piezoelectric actuators was demonstrated by [24] to develop small soft robots capable of navigating complex environments. Unfortunately, such robotic systems fail to scale due to the lack of modular structure.
@@
Once the base design of the robot is fixed, they need a set of actuators to drive the motion. These activations can be performed by external force [11, 17, 18], pressure differentials [12, 25], or by deforming the material itself [26, 27, 28, 19, 1, 10, 22]. Unfortunately, unlike natural organisms, machines cannot rely on distributed networks of active elements. Instead, their actuators must be individually connected to a control unit and a power source. Such a physical tethering limits the form of the robot and often restricts the machine to a single location in space, a glaring concern in the soft robotics community [22]. To address this issue, researchers have investigated untethered control of robotic materials [6, 7, 29, 30, 3, 31, 5, 32, 4, 33], but such methods have lacked both structural build-up of robotic surfaces using building blocks and their active control. It therefore remains an open challenge to develop an untethered actuation method that is capable of activating a large number of actuators simultaneously, required for creating modularly designed and actively controlled soft robots.
@@
Despite existing limitations, recent advances in material science and control systems have introduced new challenges for robotic design. As the number of design options increases, identifying the most optimal designs becomes increasingly complex. The computer graphics community has developed various tools to design passively controlled deformable structures, including numerical optimization models [34, 16, 26, 11, 12, 18, 10], reinforcement learning methods [35, 13, 14], and more recently, differentiable predictive models [15, 36, 37]. However, the application of these tools in the design of actively controlled robots remains limited [9]. We summarize these key aspects of robotic designs in Fig. 2, highlighting the distinctions between modular and non-modular designs, active and passive actuation in soft robots, and tethered vs untethered control systems.
@@
In this work, we seek to improve upon these long-standing challenges in soft robotics by taking inspiration from natural organisms. Instead of using specialized appendages or passively actuatable materials, we propose using an active actuator as the base material building block (see Fig. 1 a)). The robotic design is then a distributed collection of these actuators represented as edges, together forming a skeletal graph. To physically realize the robot, the building block is made from a rigid 3D-printed structure that acts as the bone. It is further equipped with two soft liquid crystal elastomer (LCE) actuators that act as muscles (see Fig. 1 b)) to shrink the bone or perform preferential bending (see Fig. 1 c)). To avoid tethering the individual muscles, we propose a novel activation system. Specifically, we modify the chemical composition of the LCE muscles to be activated by an infrared laser, thus allowing for active and untethered control of individual muscles and their parent bones. Finally, the collective behavior of the individual bones generates the final form and functionality of the robots. Our proposed material and actuation system significantly broaden the space of admissible robotic designs. To showcase our method, we first introduce three simple robots (Fig. 4): a face-bot that exhibits the basic functionalities of LCEs as muscles to control the skeleton by performing different facial animations, a $3\times 5$ grid-bot to show basic shape morphing capabilities from flat sheet to multiple target shapes in 3D and an Xbot that demonstrates untethered locomotion on flat terrain.
@@
Furthermore, to fully leverage the design possibilities that our system unlocks, we equip it with two computational tools. The first tool optimizes the skeletal design of the building blocks to create robots that can achieve a set of user-defined target configurations. To this end, we build a morph-bot that is first verified for its functionality through simulation, and then its performance is demonstrated in the real world (Fig. 5). The second tool co-optimizes the skeletal design and its control to enable the robot to perform target functions in simulation. We use this tool to design and control different frog-bots that exhibit locomotion across diverse terrains (Fig. 6). The details of the fabricated robots are summarized in Table 1. Together, these robots help us illustrate the capabilities of our universal building block of distributed actuators to build versatile robots. We believe that our computational design system for the fabrication of untethered modular robots will pave the way for future soft robotic systems.
2 Main
@@
The building block is a single line element, as shown in Figure 1. It consists of two joints and two elastic rods that form the flexible bone and two shrink actuators acting as muscles that are attached to it. We 3D-print the bones (see Sec. 5 Fabrication and Characterization of 3D-printed bones) and chemically synthesize the muscle material in the laboratory (see Sec. 5 Fabrication and Characterization of LCE-based muscles). The contraction of a single actuator bends the bone element with a designed directional bias, while the simultaneous contraction of both actuators results in the shrinkage of the bone element (see Fig. 1). We validate this experimentally in supplementary video: S1. Our building block demonstrates an excellent motion range, achieving up to 33% compression from its original length when shrinking, and a preferential curvature of ± $\pm\ 90$ degrees when bending.
Figure 3: Material Characterization: (a,b) We characterize the flexibility and bending force of our 3D-printed bones, which are influenced by the material (TPU/PLA) and the cross-section radius of the springs. (inset: 3D-printed bone with and without connectors required for shrinkage behavior of bones) (c) We demonstrate the absorption efficiency of IR820 in the infrared spectrum, showing its effectiveness in absorbing laser radiation. (d) We characterize the response by LCE muscles by varying the number of laser pulses (1-10) at the highest intensity and measure the resulting contraction, temperature changes, heating time and cooling time. Notice that as we increase the number of laser repetitions, we are able to achieve the maximum 33% contraction of our LCE muscles using the laser setup.
@@
To actuate our proposed building block, we developed a custom infrared laser heating system based on previous research [38] (see Fig. 7). We use a high-intensity near-infrared laser diode together together with a Galvoscanner. The Galvoscanner is controlled with the help of a microcontroller. To ensure precise laser actuation, we outfit our system with six RGB cameras and perform a volumetric calibration of the experimental laser area $35\times 35\ cm^{2}$. Our system transmits laser instructions in packets that encode the start and end points of the trajectory, intensity, scanning speed, and total actuation time by specifying the repetition count. We characterize the laser heating system (see Sec. 5, Designing and characterizing the control system) to ensure that it works in harmony with the untethered actuators introduced next.
@@
The actuators in our building block are rod-shaped Liquid Crystal Elastomers (LCEs). LCE muscles are heat-responsive structures that shrink in length when heated and return to their original shape when cooled to room temperature (see Fig. 1). Since the inception of LCEs in the late 1980s, the materials have attracted a significant amount of attention. Researchers have explored various ways of synthesizing and controlling LCEs [39], ranging from tethered control [40, 1] to untethered control [3, 41, 28]. However, previous methods have had limited success in incorporating untethered and active control of LCEs, resulting in restricted design space for fabricated robots. In our approach, we enable the use of LCEs as untethered muscles by incorporating an infrared-absorbing dye (IR820) into the material, allowing it to be heated via infrared radiation. This enables localized, untethered control of the LCE muscles using our proposed laser system. We characterize the LCE muscles with the laser setup (see Sec. 5 Fabrication and Characterization of LCE based muscles, and Fig. 3 (c) and (d)).
@@
The first step in creating a robot is to construct its skeletal graph ($\mathcal{N}$, $\mathcal{E}$). The skeletal graph consists of a collection of edges $\mathcal{E}$ that are physically realized with the building blocks, and the nodes $\mathcal{N}$ which represents the joints where the edges meet. The maximum number of edge elements that can be actuated by our setup to achieve target actuation is constrained by the physical limitations of the system which we explain in detail in the supplementary material.(see Appendix A: Number of actuators in a robot). From this graph structure, we automatically generate the CAD model (See Appendix A Generating the CAD file). We fabricate the model with a Fused Deposition Modeling 3D printer (see Appendix A 3D printing the skeleton). The flexibility of the built robot depends on the flexibility of the individual bones, which in turn depends on the material used and the cross-section radius. Fig. 3 (a,b) and Sec. 5: Fabrication and Characterization of 3D-printed Bones present the characterization of bones based on their material and cross-sectional radius. We chose the material and thickness for the robots based on the task at hand. We report the skeletal design details for our robots in Table 1.
@@
Our unique actuation system utilizes infrared radiation heating for untethered control of the LCE muscles and presents a particular set of challenges. Specifically, for any LCE muscle undergoing actuation, the instantaneous length varies non-linearly with temperature. This is explained by the material properties of the LCEs [42] and is experimentally reported by [43, 1]. Thus, to ensure precise actuation of the LCE muscles, we place them in precisely fabricated slots on 3D-printed bones and equip them with a tracking system. Every joint where two bones meet is equipped with an ArUco Marker that provides real-time information on the position of the bones, allowing close-loop control of the robots. (See Appendix A Tracking ArUco Markers) Now, given a task at hand, we use customized algorithms to generate laser actuation profiles that provide the laser with an instruction set to scan the individual muscles of the robot.
For shape morphing tasks, we generate a laser actuation profile that gradually transforms the flat-fabricated robot into its target shape. Since both the flat-fabricated shape and the target shape share the same connectivity, their edges form a bijection. This allows morphing between the two, given the shrinkage and curvature information for both shapes [44]. The shrinkage ratio is a continuous parameter space that we navigate by heating the muscles gradually until they reach their target contracted length. Curvature dictates the discrete choice of which muscle to heat in a bone. A positive curvature is achieved by heating the bottom muscle, while a negative curvature is achieved by heating the top muscle. The edges are then placed in a priority queue, with priority set based on how far an edge is from the target contraction ratio. To avoid overheating any given muscle, we penalize a muscle’s priority by checking the number of visits to that muscle. This heating process results in the gradual actuation of the robot until each edge reaches its target length and curvature, thus yielding the final shape.
For locomotion tasks we generate laser actuation profiles that activate the muscles, resulting in their periodic motion. The amplitude, phase, and frequency of the periodic motion are tuned for each muscle by controlling the laser intensity, scan speed, and actuation time using real-time feedback on their position.
3 Physical Robots
@@
Table 1: Table showcase the skeletal graph, their physical parameters and the intended tasks of the fabricated robots
@@
The activation of individual muscles is a basic functionality that natural organisms display. Each muscle is part of the distributed actuation system and can be controlled independently, with their combined effect results in the final shape that the organisms acquires. We demonstrate this basic functionality using the face-bot (see Fig. 4 and supplementary video S2). face-bot ’s skeletal graph consists of 22 nodes and 19 bones. Each bone is equipped with one muscle placed on the bottom side. The skeleton is 3D-printed with flexible TPU material with a 2mm2\ $m$ cross-sectional radius. We design different face animation patterns by contracting different sets of muscles. The target lengths of the animated shapes are then fed as input to the shape-morphing control algorithm (see Sec. 2 control algorithms) to generate the required laser-heating profiles. Finally, we actuate the face-bot with the laser heating profiles to obtain the desired facial expressions.
@@
The grid-bot (see Fig. 4 and supplementary video: S3) is a $3\times 5$ grid-shaped robot with its skeletal graph consisting of 15 nodes and 22 edges. The skeleton is made of flexible TPU material with a cross-sectional radius of 2mm2\ $m$, and each bone is equipped with two muscles to allow preferential bending. We provide different target configurations as input and compute the desired length and curvature per edge. We use the shape-morphing control strategy to obtain the laser-actuation profiles and apply them to morph the grid-bot into corresponding three-dimensional shapes.
Figure 5: We demonstrate how our computational design tool is used to create shape morphing robots. (a) Starting with user-provided input shapes (a) dome and saddle we use the computational design algorithm to generate the skeletal graph. This graph tessellates the input shapes into a coarse representation (b) and uses the bijection between the edges to compute the the shrinkage ratio and bending direction. We use the per-edge information to generate the laser actuation profile using the shape-morphing control algorithm. To verify the design, we use a numerical simulator (c) before fabricating the CAD model (d). Once satisfied, we 3D print the robot and attach the LCE muscles (e). Finally, we use the laser actuation profile to morph the fabricated morph-bot into the target 3D shapes (f).
@@
The range of shapes that we can achieve by morphing grid-like robots is inherently limited. The actuation profiles can fail to actuate the robot to its target configuration when gravitational forces are comparable to the robot’s bending energy. Moreover, it remains unclear how to design a grid-like skeletal graph that can closely approximate complex user-defined target configurations. To this end, we develop a tool (see Sec. 5 Computational design tool for shape morphing) that considers the fabrication constraints of our framework and generates a skeletal graph that is optimized to approximate the user-defined target configuration when actuated using untethered laser control. Furthermore, we numerically verify the functionality of both the skeletal graph and the laser actuation profile through physics-based simulation to ensure the correct behavior of the robot after fabrication (see Sec. 5 numerical simulation to verify shape-morphing robots). We use this tool to design the morph-bot .
Morph-bot is provided with a Dome and a Saddle as target shapes, and the design tool generates its skeletal graph that contains 15 nodes and 33 edges. We use this skeletal graph to first generate the CAD model of morph-bot . This CAD model is used to numerically verify morph-bot with the simulation tool. Finally, we fabricate the morph-bot using flexible TPU material with a 2mm2\ $m$ cross-section radius and equip each bone with two muscles to allow for preferential bending. We use the shape-morphing control algorithm, to morph the flat-fabricated structure to its target configurations. Fig. 5 and supplementary video: S5 show the morph-bot in action and provide physical validation of the computational design and simulation pipeline.

### v2

@@ Abstract
Natural organisms utilize distributed actuation through their musculoskeletal systems to adapt their gait for traversing diverse terrains or to morph their bodies for varied tasks. A longstanding challenge in robotics is to emulate this capability of natural organisms, which has motivated the development of numerous soft robotic systems. However, such systems are generally optimized for a single functionality, lack the ability to change form or function on demand, or remain tethered to bulky control systems. To address these limitations, we present a framework for designing and controlling robots that utilize distributed actuation. We propose a novel building block that integrates 3D-printed bones with liquid crystal elastomer (LCE) muscles as lightweight actuators, enabling the modular assembly of musculoskeletal robots. We developed LCE rods that contract in response to infrared radiation, thereby providing localized, untethered control over the distributed skeletal network and producing global deformations of the robot. To fully capitalize on the extensive design space, we introduce two computational tools: one for optimizing the robot’s skeletal graph to achieve multiple target deformations, and another for co-optimizing skeletal designs and control gaits to realize desired locomotion. We validate our framework by constructing several robots that demonstrate complex shape morphing, diverse control schemes, and environmental adaptability. Our system integrates advances in modular material building, untethered and distributed control, and computational design to introduce a new generation of robots that brings us closer to the capabilities of living organisms.
Summary
Inspired by vertebrates, we present active and untethered control of modular musculoskeletal robots optimized with computational tools.
Introduction
Figure 1: Our untethered laser heating setup (A) allows for remote and active actuation of robots. The robotic designs and controls are optimized for different functionalities (B). For every task, an optimized skeletal graph (C) is used for modular fabrication of the robot (D) with the help of base building blocks (E). Each building block is fabricated with a 3D-printed bone and equipped with two LCE muscles (F) that serve as untethered actuators. The building block can bend preferentially (G), (H) and shrink (I), depending on the selective activation of the muscle.
@@ Introduction
Natural organisms have developed remarkable mobility to traverse challenging terrains and perform complex tasks. Their ability to readily adapt to their environment lies in their structural design; instead of specialized appendages, they leverage general building blocks shaped by evolution. Vertebrates demonstrate this remarkable diversity, with different species having evolutionarily optimized skeletons tailored to their specific habitats. Further enhancing this design, a distributed network of muscles empowers these skeletons for intricate motions. This synergy creates a powerful and versatile blueprint that enables vertebrates to thrive in a wide range of environments and tasks.
@@
Figure 2: Active control of tethered soft robots (A,B)  (?, ?) has showcased rich performance for shape morphing and locomotion tasks. In contrast, untethered control achieved through different passive actuation (C)  (?) remains limited in performance. Untethered robots with active control combine the best of both approaches and allow for wireless locomotion (D)  (?). Unfortunately, such robotic systems are designed by hand and fail to scale up. From a design perspective, modular robots have shown promise by combining basic building blocks to perform shape-morphing or locomotion tasks (E)  (?). Furthermore, computational design tools allow for structural optimization of different shape morphing structures that are passively controlled (F)  (?). Modular robots created using computational design tools promise significant potential as demonstrated by the conceptual robots in (G)  (?). Our system uniquely integrates these different domains, enabling computational design of modular robots with active and untethered control to perform both shape morphing and locomotion tasks.
@@
This contrasts with the capabilities of human-designed robots that mimic natural organisms — soft robots. Soft robots have gained traction in the robotics community due to their ability to interact safely with humans or delicate environments. They are typically made of soft elastic material, where a single spatial location can generally undergo two modes of deformation: it can either bend or shrink. These modes of deformation have inspired many researchers to assemble individual blocks into structures that can be passively actuated into target shapes.  (?, ?) introduced an auxetic building block with a negative Poisson’s ratio that passively transforms a flat-fabricated configuration into target 3D shapes.  (?, ?) utilized building blocks composed of elastic beams, rigid plates, and hinge joints to design mechanically deployable structures.  (?) introduced inflatable fusing curves as building blocks in which the flat design is encoded to passively inflate to a target 3D shape.  (?, ?) used shrinkable materials as building blocks that are passively actuated with hot water to achieve the desired 3D shapes. Similar building blocks were used by  (?) to print a heat-shrinkable base sheet that self-folds into target origami objects. However, without active control of each individual unit, these systems remain limited in their ability to support general-purpose soft robotic applications.
@@
On the flip side, researchers proposed robotic materials embedded with actively controlled actuators.  (?, ?) designed robotic surfaces with high spatio-temporal control to perform shape morphing and object manipulation.  (?, ?, ?) designed shape morphing robots capable of performing locomotion tasks across a variety of terrains. Active control of piezoelectric actuators was demonstrated by  (?) to develop small soft robots capable of navigating complex environments. Unfortunately, these robotic systems are manually designed, making it difficult to generalize their framework to broader applications.
@@
Once the base design of the robot is fixed, a set of actuators is required to drive its motion. These actuators can be activated by external force  (?, ?, ?), pressure differentials  (?, ?), or by deforming the material itself  (?, ?, ?, ?, ?, ?, ?). Unfortunately, unlike natural organisms, machines cannot rely on distributed networks of active elements. Instead, their actuators must be individually connected to a control unit and a power source. Such a physical tethering limits the form of the robot and often restricts the machine to a single location in space, a glaring concern in the soft robotics community  (?). To address this issue, researchers have investigated untethered control of robotic materials  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?). However, such methods have lacked both the structural build-up of robotic surfaces using building blocks and their active control. It therefore remains an open challenge to develop an untethered actuation method that is capable of activating a large number of actuators simultaneously, required for creating modularly designed and actively controlled soft robots.
@@
Despite existing limitations, recent advances in material science and control systems have introduced new challenges for robotic design. As the number of design options increases, identifying the most optimal designs becomes increasingly complex. The computer graphics community has developed various tools to design passively controlled deformable structures, including numerical optimization models  (?, ?, ?, ?, ?, ?, ?), reinforcement learning methods  (?, ?, ?), and more recently, differentiable predictive models  (?, ?, ?). However, the application of these tools in the design of actively controlled robots remains limited  (?). We summarize these key aspects of robotic designs in Figure 2, highlighting the distinctions between modular and non-modular designs, active and passive actuation in soft robots, and tethered versus untethered control systems.
@@
In this work, we seek to address these long-standing challenges in soft robotics by taking inspiration from natural organisms. Instead of using specialized appendages or passively actuatable materials, we propose using an active actuator as the base building block (see Figure 1 (E)). The robotic design is then a distributed collection of these actuators represented as edges, together forming a skeletal graph. To physically realize the robot, the building block is made from a rigid 3D-printed structure that acts as the bone. It is further equipped with two soft liquid crystal elastomer (LCE) actuators that act as muscles (see Figure 1 (F)) to shrink the bone or perform preferential bending (see Figure 1 (G,H,I)). To avoid tethering the individual muscles, we propose a novel activation system. We achieve this by modifying the chemical composition of the LCE muscles so they can be activated by an infrared laser, enabling active and untethered control of individual muscles and their parent bones. Finally, the collective behavior of the individual bones generates the final form and functionality of the robots. Our proposed material and actuation system significantly broaden the space of admissible robotic designs. To showcase our method, we first introduce three simple robots (Figure 4): a face-bot that exhibits the basic functionalities of LCEs as muscles to control the skeleton by performing different facial animations, a $3\times 5$ grid-bot to show basic shape morphing capabilities from flat sheet to multiple target shapes in 3D and an Xbot that demonstrates untethered locomotion on flat terrain.
@@
Furthermore, to fully leverage the design possibilities that our system unlocks, we equip it with two computational tools. The first tool optimizes the skeletal design of the building blocks to create robots that can achieve a set of user-defined target configurations. To this end, we build a morph-bot that is first verified for its functionality through simulation, and then its performance is demonstrated in the real world (Figure 5). The second tool co-optimizes the skeletal design and its control to enable the robot to perform target functions in simulation. We use this tool to design and control different frog-bots that exhibit locomotion across diverse terrains (Figure 6). The details of the fabricated robots are summarized in Table 1. These robots collectively illustrate the potential of our universal building block of distributed actuators for creating versatile robots. We believe that our computational design system for the fabrication of untethered modular robots will pave the way for future soft robotic systems.
Results
@@
The building block is a single line element, as shown in Figure 1. It consists of two joints and two elastic rods that form the flexible bone and two shrink actuators acting as muscles that are attached to it. We 3D-print the bones (see Materials and Methods Materials and Methods: Fabrication and Characterization of 3D-printed bones) and chemically synthesize the muscle material in the laboratory (see Materials and Methods Materials and Methods: Fabrication and Characterization of LCE-based muscles). The contraction of a single actuator bends the bone element with a designed directional bias, while the simultaneous contraction of both actuators results in the shrinkage of the bone element (see Figure 1 (E)). We validate this experimentally in supplementary video: S1. Our building block demonstrates an excellent motion range, achieving up to 33% compression from its original length when shrinking, and a preferential curvature of ± $\pm\ 90$ degrees when bending.
Figure 3: Material Characterization: (A,B) We characterize the flexibility and bending force of our 3D-printed bones, which are influenced by the material (TPU/PLA) and the cross-section radius of the springs. (inset: 3D-printed bone with and without connectors required for shrinkage behavior of bones) (C) We demonstrate the absorption efficiency of IR820 in the infrared spectrum, showing its effectiveness in absorbing laser radiation (D). We characterize the response of LCE muscles by varying the number of times (1–10) the laser is swept at the highest intensity and measuring the resulting contraction, temperature changes, heating time, and cooling time. Notice that as we increase the number of laser repetitions, we are able to achieve the maximum 33% contraction of our LCE muscles using the laser setup.
@@
To actuate our proposed building block, we developed a custom infrared laser heating system based on previous research  (?) (see Figure S1). We use a high-intensity near-infrared laser diode together with a Galvoscanner. The Galvoscanner is controlled with the help of a microcontroller. To ensure precise laser actuation, we equip our system with six RGB cameras and perform volumetric calibration of the $35\times 35\ cm^{2}$ experimental laser area. Our system transmits laser instructions in packets that encode the start and end points of the trajectory, intensity, scanning speed, and total actuation time by specifying the repetition count. We characterize the laser heating system (see Designing and Characterizing the Control System Materials and Methods: Designing and characterizing the control system) to ensure that it works in harmony with the untethered actuators introduced next.
@@
The actuators in our building block are rod-shaped Liquid Crystal Elastomers (LCEs). LCE muscles are heat-responsive structures that shrink in length when heated and return to their original shape when cooled to room temperature (see Figure 1). Since the inception of LCEs in the late 1980s, the materials have attracted a significant amount of attention. Researchers have explored various ways of synthesizing and controlling LCEs  (?), ranging from tethered control  (?, ?) to untethered control  (?, ?, ?). However, previous methods have had limited success in incorporating untethered and active control of LCEs, resulting in restricted design space for fabricated robots. In our approach, we enable the use of LCEs as untethered muscles by incorporating an infrared-absorbing dye (IR820) into the material, allowing it to be heated via infrared radiation. This enables localized, untethered control of the LCE muscles using our proposed laser system. We characterize the LCE muscles with the laser setup (see Fabrication and Characterization of LCE based Muscles Materials and Methods: Fabrication and Characterization of LCE based muscles, and Figure 3 (C) and (D)).
@@
The first step in creating a robot is to construct its skeletal graph ($\mathcal{N}$, $\mathcal{E}$). The skeletal graph consists of a collection of edges $\mathcal{E}$ that are physically realized with the building blocks, and the nodes $\mathcal{N}$ which represents the joints where the edges meet. From this graph structure, we automatically generate the CAD model. We fabricate the model with a Fused Deposition Modeling 3D printer. The flexibility of the built robot depends on the flexibility of the individual bones, which in turn depends on the material used and the cross-section radius as illustrated in Figure 3 (A,B) and Fabrication and Characterization of 3D-printed Bones: Materials and Methods: Fabrication and Characterization of 3D-printed Bones. We chose the material and thickness for the robots based on the task at hand. We report the skeletal design details for our robots in Table 1. The maximum number of edge elements that can be actuated by our setup to achieve the target configuration is constrained by the physical limitations of the system. We explain these constraints, along with other design and fabrication aspects in the supplementary material (see Supplementary Information Supplementary Information: Number of actuators in a robot, Generating the CAD file, 3D printing the skeleton).
@@
Our unique actuation system utilizes infrared radiation heating for untethered control of the LCE muscles which presents a particular set of challenges. Specifically, for any LCE muscle undergoing actuation, the instantaneous length varies non-linearly with temperature. This is explained by the material properties of the LCEs  (?) and is experimentally reported by  (?, ?). Thus, to ensure precise actuation of the LCE muscles, we place them in precisely fabricated slots on 3D-printed bones and equip them with a tracking system. Every joint where two bones meet is equipped with an ArUco Marker that provides real-time information on the position of the bones, allowing close-loop control of the robots (details in Supplementary Information Supplementary Information: Tracking ArUco Markers). Now, given a task at hand, we use customized algorithms to generate laser actuation profiles that provide the laser with an instruction set to scan the individual muscles of the robot.
For shape morphing tasks, we generate a laser actuation profile that gradually transforms the flat-fabricated robot into its target shape. Since both the flat-fabricated shape and the target shape share the same connectivity, their edges form a bijection. This allows morphing between the two, given the shrinkage and curvature information for both shapes  (?). The shrinkage ratio is a continuous parameter that we navigate by heating the muscles gradually until they reach their target contracted length. Curvature dictates the discrete choice of which muscle to heat in a bone. A positive curvature is achieved by heating the bottom muscle, while a negative curvature is achieved by heating the top muscle. The edges are then placed in a priority queue, with priority set based on how far an edge is from the target contraction ratio. To avoid overheating any given muscle, we penalize a muscle’s priority by checking the number of visits to that muscle. This heating process results in the gradual actuation of the robot until each edge reaches its target length and curvature, thus yielding the final shape.
For locomotion tasks, we generate laser actuation profiles that activate the muscles, resulting in their periodic motion. The amplitude, phase, and frequency of the periodic motion are tuned for each muscle by controlling the laser intensity, scan speed, and actuation time using real-time feedback on their position.
Physical Robots
@@
Table 1: Table showcases the skeletal graphs, their physical parameters and the intended tasks of the fabricated robots
@@
Activation of individual muscles is a fundamental capability exhibited by natural organisms. Each muscle is part of the distributed actuation system and can be controlled independently, with their combined effect results in the final shape that the organisms acquires. We demonstrate this basic functionality using the face-bot (see Figure 4 and supplementary video S2). face-bot ’s skeletal graph consists of 22 nodes and 19 bones. Each bone is equipped with one muscle placed on the bottom side. The skeleton is 3D-printed with flexible TPU material with a 2mm2\ $m$ cross-sectional radius. We design different face animation patterns by contracting different sets of muscles. The target lengths of the animated shapes are then fed as input to the shape-morphing control algorithm to generate the required laser-heating profiles. Finally, we actuate the face-bot with the laser heating profiles to obtain the desired facial expressions.
@@
The grid-bot (see Figure 4 and supplementary video: S3) is a $3\times 5$ grid-shaped robot with its skeletal graph consisting of 15 nodes and 22 edges. The skeleton is made of flexible TPU material with a cross-sectional radius of 2mm2\ $m$, and each bone is equipped with two muscles to allow preferential bending. We provide different target configurations as input and compute the desired length and curvature per edge. We use the shape-morphing control strategy to obtain the laser-actuation profiles and apply them to morph the grid-bot into corresponding three-dimensional shapes.
Figure 5: We demonstrate how our computational design tool is used to create shape morphing robots. Starting with user-provided input shapes (A) dome and saddle, we use a computational design tool to generate the skeletal graph. This graph tessellates the input shapes into a coarse representation (B) and uses the bijection between the edges to compute the shrinkage ratio and bending direction. We use the per-edge information to generate the laser actuation profile with the shape-morphing control algorithm. Before fabrication, we use a numerical simulator (C) to verify the design (D). After verification, we 3D print the robot and attach the LCE muscles (E). Finally, we use the laser actuation profile to morph the fabricated morph-bot towards the target 3D shapes (F). The system achieves the dome shape with good visual accuracy but only actuates the saddle to a partial configuration where the sharp rising edges of the desired saddle shape are countered by gravitation force.
@@
The range of shapes that we can achieve by morphing grid-like robots is inherently limited. The actuation profiles can fail to actuate the robot to its target configuration when gravitational forces are comparable to the robot’s bending energy. Moreover, it remains unclear how to design a grid-like skeletal graph that can closely approximate complex user-defined target configurations. To this end, we develop a tool (see Material and Methods: Computational design tool for shape morphing) that considers the fabrication constraints of our framework and generates a skeletal graph that is optimized to approximate the user-defined target configuration when actuated using untethered laser control. Furthermore, we numerically verify the functionality of both the skeletal graph and the laser actuation profile through physics-based simulation to ensure the correct behavior of the robot after fabrication. We use this tool to design the morph-bot .
Morph-bot is provided with a Dome and a Saddle as target shapes, and the design tool generates its skeletal graph that contains 15 nodes and 33 edges. We use this skeletal graph to first generate the CAD model of morph-bot . This CAD model is used to numerically verify morph-bot with the simulation tool. Finally, we fabricate the morph-bot using flexible TPU material with a 2mm2\ $m$ cross-section radius and equip each bone with two muscles to allow for preferential bending. We use the shape-morphing control algorithm, to morph the flat-fabricated structure to its target configurations. Figure 5 and supplementary video: S5 show the morph-bot in action and provide physical validation of the computational design and simulation pipeline. The morph-bot successfully transforms into a dome shape with good visual fidelity but achieves only a partial transformation into the saddle shape. This limitation underscores the simulation-to-reality gap as the gravitational forces dominate the TPU made morph-bot , preventing the structure to reach complete transformation and presents venue for future research in customized physics based simulation.